# Inverse Hamiltonian reduction in type A and generalized slices in the affine Grassmannian

### v1

@@ Abstract.
We give a geometric proof of inverse Hamiltonian reduction for all finite W-algebras in type $A$, a certain embedding of the finite W-algebra corresponding to an arbitrary nilpotent in $\mathfrak{gl}_{N}$ into that corresponding to a larger nilpotent with respect to the closure order on orbits, tensored with an auxiliary algebra of differential operators. We first prove a classical analogue for equivariant Slodowy slices using multiplication maps on generalized slices in the affine Grassmannian, then deduce the result for equivariant finite W-algebras by Fedosov quantization. This implies the statement for finite W-algebras, as well as Kostant-Whittaker reductions of arbitrary algebras in the category of Harish Chandra bimodules, including quantizations of Moore–Tachikawa varieties.
@@ 1. Introduction
The finite W-algebra ${\mathcal{W}}_{f}({{\mathfrak{g}}})$ associated to a nilpotent element $f$ in a semisimple Lie algebra ${{\mathfrak{g}}}$ is an associative algebra quantizing the transverse slice $S_{f}\subset{{\mathfrak{g}}}$ to the nilpotent orbit $\mathbb{O}_{f}=G\cdot f\subset{{\mathfrak{g}}}$. In type $A_{N-1}$, such orbits are parameterized by partitions of $N$, or equivalently tuples of positive integers $\mu=(\mu_{1}\geq...\geq\mu_{n-1}\geq 0)$ such that the total sum $|\mu|=\sum_{k=1}^{n-1}\mu_{k}=N$; we call the positive integer $n-1$ the length of such a partition, and identify the set of partitions of length at most $n-1$ with the set of dominant, integral coweights $\Lambda^{+}({{\textup{PGL}}}_{n})$ for the projective linear group ${{\textup{PGL}}}_{n}$.
In the case $\mu=[1^{N}]=(1,...,1)$ for which $n-1=N$, the corresponding nilpotent $f_{\mu}=0\in{{\mathfrak{gl}}}_{N}$, so that the slice $S_{0}={{\mathfrak{g}}}$ and the finite W-algebra ${\mathcal{W}}_{0}({{\mathfrak{g}}})=U({{\mathfrak{g}}})$ is simply the universal enveloping algebra of ${{\mathfrak{g}}}$, which naturally controls the representation theory of ${{\mathfrak{g}}}$. The algebras ${\mathcal{W}}_{f_{\mu}}({{\mathfrak{g}}})$ are a natural family of generalizations thereof, which analogously control the categories of so-called Whittaker modules for ${{\mathfrak{g}}}$, with respect to a unipotent group $N_{\mu}$ and character $\chi_{\mu}=\langle f_{\mu},\cdot\rangle_{{\mathfrak{g}}}\in\mathfrak{n}_{\mu}^{\vee}$ determined by $f_{\mu}$ (up to some relatively insignificant auxiliary choices).
@@
There exists a isomorphism of Poisson algebras with Hamiltonian $\mathbb{G}_{\alpha}$-action
@@
[…]
@@
In particular, we deduce the following corollary for the quantized coordinate rings of Moore–Tachikawa varieties:
@@ 1.1. Relation with previous results
Though perhaps less obviously than the other results we discuss below, Theorems 1.2 and 1.1 can be understood as generalizations of the main results of [KPW22], and moreover our proof of the former follows from first proving a direct generalization of loc. cit. For any $n\in\mathbb{N}$, dominant coweight $\lambda\in\Lambda^{+}_{{{\textup{GL}}}_{n}}$, and arbitrary coweight $\mu\in\Lambda_{{{\textup{PGL}}}_{n}}$, we let $\mathrm{Gr}_{\mu}^{\lambda}$ denote the generalized slice between orbits in the affine Grassmannian of ${{\textup{PGL}}}_{n}$, and we prove:
@@
We then use certain isomorphisms relating equivariant Slodowy slices in type $A$ with generalized slices in the affine Grassmannian of ${\rm PGL}_{2N}$. Theorem 1.2 is therefore a special case of this more general result on affine Grassmannian slices.
@@
The results of this paper can also be understood as a generalization of the main theorem of [KPW22] to the case of higher rank additive group schemes $\mathbb{G}_{\alpha}$, translated under the so-called Mirković–Vybornov isomorphism [MVK22] to the setting of slices to nilpotent orbits and in turn finite W-algebras. Indeed, our proof of Theorem 1.2, on which the remainder of our results are built, follows the proof in [KPW22] very closely. Moreover, the results of the companion papers [BN25a] and [BN25b] should be understood as part of a broader program to extend the results of [FKP+18] and [WWY20], and in turn [BK04b], to the affine setting in a compatible way.
@@
This result was recently re-popularized by Dražen Adamović, who showed in [Ada04, Ada19] that the result above in type $A_{1}$ descends to an embedding of the simple quotient vertex algebras, and used this to relate their categories of modules in order to express that category of modules over the affine algebra in terms of the more simple category of modules over the principal affine W-algebra for $\mathfrak{sl}_{2}$. Following this, many papers were written studying this phenomena in increasingly general families of examples: see [ACG24], [Feh23b], [Feh23a], [CFLN24], [FFFN24], and [FKN25]. The proof of the analogous embeddings in all of these papers rely on analogous ad hoc combinatorial arguments relating the explicit descriptions of the vertex algebras in terms of so-called screening operators. The goal of this series of papers is to establish the general result in type $A$ at generic level $\kappa$ and for arbitrary nilpotents, as well as to give a geometric proof providing a more conceptual explanation for the results.
Finally, both this work and its companion should be viewed as part of a series of recent works in the vertex algebra literature where free field realizations are constructed geometrically—via localization. In particular, the viewpoint that (sheaves of) vertex algebras should be viewed as geometric objects over a space and that free field realizations should be viewed as arising from restricting these sheaves to open subsets. This has been used to great effect in [ACM11], [BMR19], [BN23], [But23], [BR23], [Fur23], [AKM23], [BDM+24].
1.2. Acknowledgments
The authors would like to thank Dražen Adamović, Tomoyuki Arakawa, Justine Fasquel, Zachary Fehily, Naoki Genra, Thibault Juillard, Joel Kamnitzer, Vasily Krylov, Shigenori Nakatsuka, Alex Weekes, and Yehao Zhou for useful discussions. We would like to especially thank Christopher Beem, whose collaboration in the early stages of this project was invaluable.
@@ 2.1. The affine Grassmannian
Let $G$ be an affine algebraic group, $X$ a smooth algebraic curve over $\mathbb{\mathbb{C}}$, and for a closed point $x\in X$ let $\mathbb{D}_{x}=X^{\wedge}_{\{x\}}={\textup{Spf}\ }\mathcal{O}_{x}$ be the formal disk at $x$ and $\mathbb{D}_{x}^{\times}=X^{\wedge}_{\{x\}}-\{x\}={\textup{Spf}\ }\mathcal{K}_{x}$, the punctured formal disk, where $\mathcal{O}_{x}=\mathcal{O}_{X,x}^{\wedge}$ denotes the complete local ring at $x\in X$ and $\mathcal{K}_{x}$ denotes its field of fractions. We also write simply $\mathcal{O}=\mathbb{\mathbb{C}}{{[\![t]\!]}}$ and $\mathcal{K}=\mathbb{\mathbb{C}}{{(\!(t)\!)}}$ as well as $\mathbb{D}=(\mathbb{A}^{1})^{\wedge}_{\{0\}}={\textup{Spf}\ }\mathcal{O}$ and $\mathbb{D}^{\times}=(\mathbb{A}^{1})^{\wedge}_{\{0\}}\setminus\{0\}={\textup{Spf}\ }\mathcal{K}$ for the model examples on $\mathbb{A}^{1}$ at 00, which are (non-canonically) isomorphic to those at any smooth point $x$ on any curve $X$, by the Cohen structure theorem.
@@
$G(\mathcal{K})=\bigsqcup_{\lambda\in\Lambda^{+}_{G}}G(\mathcal{O})t^{\lambda}G(\mathcal{O})\quad\quad\text{and thus}\quad\quad\mathrm{Gr}_{G}=\bigsqcup_{\lambda\in\Lambda_{G}^{+}}\mathrm{Gr}_{G}^{\lambda}\quad\quad\text{where}\quad\quad\mathrm{Gr}_{G}^{\lambda}=G(\mathcal{O})t^{\lambda}$
@@
(2.1.2) $\operatorname{Bun}_{G}(X)\xrightarrow{\cong}\left[G(\mathcal{O}_{X\setminus\{x\}})\backslash G(\mathcal{K}_{x})/G(\mathcal{O}_{x})\right]=\left[G(\mathcal{O}_{X\setminus\{x\}})\backslash\mathrm{Gr}_{G,x}\right]$
@@
(2.1.4) $\operatorname{Bun}_{G}(\mathbb{P}^{1})\xrightarrow{\cong}\left[G[t^{-1}]\backslash G[t^{\pm 1}]/G[t]\right]\quad\quad\text{and thus}\quad\quad\mathrm{Gr}_{G}\cong\left[G[t^{\pm 1}]/G[t]\right]\ .$
@@ 2.2. Generalized slices in the affine Grassmannian
$\operatorname{Bun}_{G}(\mathbb{P}^{1})_{\infty}\xrightarrow{\cong}\left[G_{1}{{[\![t^{-1}]\!]}}\backslash G{{(\!(t^{-1})\!)}}/G[t]\right]\quad\quad\text{and}\quad\quad\operatorname{Bun}_{G}(\mathbb{P}^{1})_{(B,\infty)}\xrightarrow{\cong}\left[I_{\infty}\backslash G{{(\!(t^{-1})\!)}}/G[t]\right]$
@@
[…]
@@
$\operatorname{Bun}_{B}(\mathbb{P}^{1})\to\operatorname{Bun}_{G}(\mathbb{P}^{1})_{(B,\infty)}\quad\quad\text{modeled by}\quad\quad\left[B{{[\![t^{-1}]\!]}}\backslash B{{(\!(t^{-1})\!)}}/B[t]\right]\to\left[I_{\infty}\backslash G{{(\!(t^{-1})\!)}}/G[t]\right]\ ,$
@@
  • $\sigma:\mathcal{P}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}^{\textup{triv}}$ is a trivialization of the restriction $\mathcal{P}|_{\mathbb{A}^{1}_{\infty}}\in\operatorname{Bun}_{G}(\mathbb{A}^{1}_{\infty})$, with pole at 00 of degree bounded by $\lambda$;
@@
[…]
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(X)$ is a principal $G$-bundle on $X$ for $i=1,...,n$,
@@
  • $\sigma_{i}:\mathcal{P}_{i}|_{X\setminus\{x\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{X\setminus\{x\}}$ is an isomorphism of principal $G$-bundles for $i=2,...,n$.
@@
where we let $G(\mathcal{K})\times_{G(\mathcal{O})}G(\mathcal{K})=[G(\mathcal{K})/G(\mathcal{O})]\times_{[{\textup{pt}}/G(\mathcal{O})]}[G(\mathcal{O})\backslash G(\mathcal{K})]=\left[(G(\mathcal{K})\times G(\mathcal{K}))/G(\mathcal{O})\right]$, the diagonal quotient by the product of the $G(\mathcal{O})$-actions on the right on the first factor and on the left on the second, and similarly the further iterated products are defined inductively. Note there is a natural convolution map
@@
For $\underline{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ an $n$-tuple of dominant coweights with sum $\lambda$, we have the natural substack $\overline{\mathrm{Gr}}^{\underline{\lambda}}_{G}\subset\mathrm{Gr}^{(n)}_{G}$ defined by the condition that each map $\sigma_{i}$ has a pole at 00 bounded by $\lambda_{i}$, or equivalently in terms of the quotient stack description
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $\mathbb{P}^{1}$ for $i=1,...,n$;
  • $\sigma_{1}:\mathcal{P}_{1}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}^{\textup{triv}}$ is a trivialization with pole at 00 bounded by $\lambda_{1}$;
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{A}^{1}_{\infty}}$ is an isomorphism of principal $G$-bundles with pole at 00 bounded by $\lambda_{i}$ for $i=2,...,n$;
@@ 2.3. Beilinson–Drinfeld generalized slices
[…]
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(X)$ is a principal $G$-bundle on $X$ for $i=1,...,n$;
@@
  • $\sigma_{i}:\mathcal{P}_{i}|_{X\setminus\{x_{i}\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{X\setminus\{x_{i}\}}$ is an isomorphism of principal $G$-bundles for $i=2,...,n$.
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $X$ for $i=1,...,n$;
@@
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}$ is an isomorphism of principal $G$-bundles with pole at $x_{i}$ bounded by $\lambda_{i}$ for $i=2,...,n$;

### v2

@@ Abstract.
We give a geometric proof of inverse Hamiltonian reduction for all finite W-algebras in type $A$, a certain embedding of the finite W-algebra corresponding to an arbitrary nilpotent in $\mathfrak{gl}_{N}$ into that corresponding to a larger nilpotent with respect to the closure order on orbits, tensored with an auxiliary algebra of differential operators. We first prove a classical analogue for equivariant Slodowy slices using multiplication maps on generalized slices in the affine Grassmannian, then deduce the result for equivariant finite W-algebras by Fedosov quantization. This implies the statement for finite W-algebras, as well as Kostant–Whittaker reductions of arbitrary algebras in the category of Harish Chandra bimodules, including quantizations of Moore–Tachikawa varieties.
@@ 1. Introduction
The finite W-algebra ${\mathcal{W}}_{f}({{\mathfrak{g}}})$ associated to a nilpotent element $f$ in a semisimple Lie algebra $\mathfrak{g}$}} is an associative algebra quantizing the transverse slice $S_{f}\subset{{\mathfrak{g}}}$ to the nilpotent orbit $\mathbb{O}_{f}=G\cdot f\subset{{\mathfrak{g}}}$. In type $A_{N-1}$, such orbits are parameterized by partitions of $N$, or equivalently tuples of positive integers $\mu=(\mu_{1}\geq...\geq\mu_{n-1}\geq 0)$ such that the total sum $|\mu|=\sum_{k=1}^{n-1}\mu_{k}=N$; we call the positive integer n1n-1 the length of such a partition, and identify the set of partitions of length at most n1n-1 with the set of dominant, integral coweights $\Lambda^{+}({{\textup{PGL}}}_{n})$ for the projective linear group ${{\textup{PGL}}}_{n}$.
In the case $\mu=[1^{N}]=(1,...,1)$ for which n1=Nn-1=N, the corresponding nilpotent $f_{\mu}=0\in{{\mathfrak{gl}}}_{N}$, so that the slice $S_{0}={{\mathfrak{g}}}$ and the finite W-algebra ${\mathcal{W}}_{0}({{\mathfrak{g}}})=U({{\mathfrak{g}}})$ is simply the universal enveloping algebra of $\mathfrak{g}$}}, which naturally controls the representation theory of $\mathfrak{g}$}}. The algebras ${\mathcal{W}}_{f_{\mu}}({{\mathfrak{g}}})$ are a natural family of generalizations thereof, which analogously control the categories of so-called Whittaker modules for $\mathfrak{g}$}}, with respect to a unipotent group $N_{\mu}$ and character $\chi_{\mu}=\langle f_{\mu},\cdot\rangle_{{\mathfrak{g}}}\in\mathfrak{n}_{\mu}^{\vee}$ determined by $f_{\mu}$ (up to some relatively insignificant auxiliary choices).
@@
There exists an isomorphism of Poisson algebras with Hamiltonian $\mathbb{G}_{\alpha}$-action
@@
${\mathcal{W}}_{\hbar,f_{\mu}}({{\mathfrak{gl}}}_{N})=\Gamma(S_{f_{\mu}},\mathcal{A}_{f_{\mu}})\xrightarrow{\text{res}}\Gamma(\mathbb{G}_{\alpha}\times U\times S_{f_{\mu+\alpha}},\mathcal{D}^{\rm loc}_{\alpha}\boxtimes\mathcal{A}_{f_{\mu+\alpha}})=D_{\hbar,\alpha}^{\rm loc}\otimes{\mathcal{W}}_{\hbar,f_{\mu+\alpha}}({{\mathfrak{gl}}}_{N})\ .$
@@
In particular, we deduce the following corollary for the natural quantizations of coordinate rings of non-principal Moore–Tachikawa varieties in type A:
@@ 1.1. Relation with previous results
Though perhaps less obviously than the other results we discuss below, Theorems 1.2 and 1.1 can be understood as generalizations of the main results of [KPW22], and moreover our proof of the former follows from first proving a direct generalization of loc. cit. For any $n\in\mathbb{N}$, dominant coweight $\lambda\in\Lambda^{+}_{{{\textup{PGL}}}_{n}}$, and arbitrary coweight $\mu\in\Lambda_{{{\textup{PGL}}}_{n}}$, we let $\mathrm{Gr}_{\mu}^{\lambda}$ denote the generalized slice between orbits in the affine Grassmannian of ${{\textup{PGL}}}_{n}$, and we prove:
@@
We then use certain isomorphisms relating equivariant Slodowy slices in type $A$ with generalized slices in the affine Grassmannian of ${\rm PGL}_{2N}$, extending the Mirković–Vybornov isomorphism [MVK22], to deduce Theorem 1.2 above from the preceding result. Moreover, the results of the companion papers [BN25a] and [BN25b] can be understood as part of a broader program to extend the results of [FKP+18] and [WWY20], and in turn [BK04b], to the setting of affine Lie algebras, some aspects of which were conjectured in [BR23].
@@
This result was recently re-popularized by Dražen Adamović, who showed in [Ada05, Ada19] that the result above in type $A_{1}$ descends to an embedding of the simple quotient vertex algebras, and used this to relate their categories of modules in order to express that category of modules over the affine algebra in terms of the more simple category of modules over the principal affine W-algebra for $\mathfrak{sl}_{2}$. Following this, many papers were written studying this phenomena in increasingly general families of examples: see [ACG24], [Feh23b], [Feh23a], [CFLN24], [FFFN24], and [FKN25]. The proof of the analogous embeddings in all of these papers rely on analogous ad hoc combinatorial arguments relating the explicit descriptions of the vertex algebras in terms of so-called screening operators. The goal of this series of papers is to establish the general result in type $A$ at generic level $\kappa$ and for arbitrary nilpotents, as well as to give a geometric proof providing a more conceptual explanation for the results.
Finally, both this work and its companion should be viewed as part of a series of recent works in the vertex algebra literature where free field realizations are constructed geometrically—via localization. In particular, the viewpoint that (sheaves of) vertex algebras should be viewed as geometric objects over a space and that free field realizations should be viewed as arising from restricting these sheaves to open subsets. This has been used to great effect in [ACM11], [BMR19], [BN23], [BF23], [But23], [Fur23], [AKM23], [BDM+24].
Acknowledgments
The authors would like to thank Dražen Adamović, Tomoyuki Arakawa, Justine Fasquel, Zachary Fehily, Naoki Genra, Thibault Juillard, Joel Kamnitzer, Vasily Krylov, Shigenori Nakatsuka, Alex Weekes, and Yehao Zhou for useful discussions. We would like to especially thank Christopher Beem, whose collaboration in the early stages of this project was invaluable. The authors would also like to thank Ivan Losev for bringing to our attention an error in an earlier draft of this work.
We gratefully acknowledge the support of ERC grant # 864828, and for D.B. the support of the Simons Collaboration - New Structures in Low-dimensional Topology grant.
@@ 2.1. The affine Grassmannian
Let $G$ be an affine algebraic group, $X$ a smooth algebraic curve over $\mathbb{\mathbb{C}}$, and for a closed point $x\in X$ let $\mathbb{D}_{x}=X^{\wedge}_{\{x\}}={\textup{Spf}\ }\mathcal{O}_{x}$ be the formal disk at $x$ and $\mathbb{D}_{x}^{\times}=X^{\wedge}_{\{x\}}-\{x\}={\textup{Spf}\ }\mathcal{K}_{x}$, the punctured formal disk, where $\mathcal{O}_{x}=\mathcal{O}_{X,x}^{\wedge}$ denotes the complete local ring at $x\in X$ and $\mathcal{K}_{x}$ denotes its field of fractions. We also write simply $\mathcal{O}=\mathbb{\mathbb{C}}{{[\![t]\!]}}$ and $\mathcal{K}=\mathbb{\mathbb{C}}{{(\!(t)\!)}}$ as well as $\mathbb{D}=(\mathbb{A}^{1})^{\wedge}_{\{0\}}={\textup{Spf}\ }\mathcal{O}$ and $\mathbb{D}^{\times}=(\mathbb{A}^{1})^{\wedge}_{\{0\}}\setminus\{0\}={\textup{Spf}\ }\mathcal{K}$ for the model examples on $\mathbb{A}^{1}$ at 0, which are (non-canonically) isomorphic to those at any smooth point $x$ on any curve $X$, by the Cohen structure theorem.
@@
[…]
@@
[…]
@@
[…]
@@ 2.2. Generalized slices in the affine Grassmannian
[…]
@@
defined by restricting the trivialization from $\mathbb{D}_{\infty}$ to $\{\infty\}$, forgetting the trivialization at $\{\infty\}$ to a choice of $B$-structure, and forgetting the $B$-structure entirely, respectively.
@@
[…]
@@
  • $\sigma:\mathcal{P}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}^{\textup{triv}}$ is a trivialization of the restriction $\mathcal{P}|_{\mathbb{A}^{1}_{\infty}}\in\operatorname{Bun}_{G}(\mathbb{A}^{1}_{\infty})$, with pole at 0 of degree bounded by $\lambda$;
@@
    $\mathcal{P}_{B}|_{\mathbb{A}^{1}_{\infty}}\to\mathcal{P}_{B}|_{\mathbb{A}^{1}_{\infty}}\times_{B}G\xrightarrow{\phi|_{\mathbb{A}^{1}_{\infty}}}\mathcal{P}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\sigma}\mathcal{P}^{\textup{triv}}=\mathbb{A}^{1}_{\infty}\times G$
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(X)$ is a principal $G$-bundle on $X$ for i=1,,ni=1,...,n,
@@
  • $\sigma_{i}:\mathcal{P}_{i}|_{X\setminus\{x\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{X\setminus\{x\}}$ is an isomorphism of principal $G$-bundles for i=2,,ni=2,...,n.
@@
[…]
@@
For $\underline{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ an $n$-tuple of dominant coweights with sum $\lambda$, we have the natural substack $\overline{\mathrm{Gr}}^{\underline{\lambda}}_{G}\subset\mathrm{Gr}^{(n)}_{G}$ defined by the condition that each map $\sigma_{i}$ has a pole at 0 bounded by $\lambda_{i}$, or equivalently in terms of the quotient stack description
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $\mathbb{P}^{1}$ for i=1,,ni=1,...,n;
  • $\sigma_{1}:\mathcal{P}_{1}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}^{\textup{triv}}$ is a trivialization with pole at 0 bounded by $\lambda_{1}$;
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{A}^{1}_{\infty}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{A}^{1}_{\infty}}$ is an isomorphism of principal $G$-bundles with pole at 0 bounded by $\lambda_{i}$ for i=2,,ni=2,...,n;
@@ 2.3. Beilinson–Drinfeld generalized slices
    $\mathcal{P}_{B}|_{\mathbb{D}_{\infty}}\to\mathcal{P}_{B}|_{\mathbb{D}_{\infty}}\times_{B}G\xrightarrow{\phi|_{\mathbb{D}_{\infty}}}\mathcal{P}|_{\mathbb{D}_{\infty}}\xrightarrow{\sigma}\mathcal{P}^{\textup{triv}}=\mathbb{D}_{\infty}\times G$
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(X)$ is a principal $G$-bundle on $X$ for i=1,,ni=1,...,n;
@@
  • $\sigma_{i}:\mathcal{P}_{i}|_{X\setminus\{x_{i}\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{X\setminus\{x_{i}\}}$ is an isomorphism of principal $G$-bundles for i=2,,ni=2,...,n.
@@
  • $\mathcal{P}_{i}\in\operatorname{Bun}_{G}(\mathbb{P}^{1})$ is a principal $G$-bundle on $X$ for i=1,,ni=1,...,n;
@@
  • $\sigma_{i}:\mathcal{P}_{i}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}\xrightarrow{\cong}\mathcal{P}_{i-1}|_{\mathbb{P}^{1}\setminus\{x_{i}\}}$ is an isomorphism of principal $G$-bundles with pole at $x_{i}$ bounded by $\lambda_{i}$ for i=2,,ni=2,...,n;